LHC 2012 - latest news

Week 31: Machine coordinators: Gianluigi Arduini, Eva Barbara Holzer

Plan for the week: Luminosity Production

Access planned around 15:00 Monday 30th July for VIP visit

Date

Summary

Link to report from morning meeting

Monday 6 August
  • 1:53 start injecting fill 2915, average bunch intensity 1.53E11, good ramp and squeeze
  • 3:13 dump on instability 70 sec after the collision beam process was launched due to RS9 losses on TCTH.4R5.B2
  • 3:52 changing the thyristor did not solve the problem with the ALICE solenoid is staying off
  • 4:20 SPS extraction kicker problem
8:30 meeting
Sunday 5 August
  • 05:23 cryo cold compressor down - access
  • also access for ALICE - afterwards trip of ALICE solenoid during ramping up
  • 21:51 cryo recovered after 16.5 hours, start pre-cycle, ALICE solenoid still in fault (broken thyristor)
9:00 meeting
Saturday 4 August
  • 0:45 dump during filling: ALICE dipole tripped (piquet checking)
  • 1:30 ALICE dipole re-set, injecting again
  • Injection with additional empty LHC cycle in the PSB/PS in front of the 4 LHC cycles. Improvement in the uniformity of the emittance of the trains.
  • 1:56 dump during filling: LBDS self-trigger (MKD B2 kicker in fault)
  • Access needed: AUE reset by piquet
  • 4:00 pre-cycle
  • 4:41 SPS main power converters down (earth fault)
  • 5:11 CMS detector partly down
  • 5:50 SPS back and CMS ready for injection – re-injecting for fill 2911. Average bunch intensity 1.53E11 at injection. Low losses during the ramp but instability developing at flat-top while waiting for solution of controls problem.
  • Heavy losses on B2 during and at the end of squeeze. Fill saved by chromaticity trim (+5 H/V B1 and +3 H/V B2)
  • 7:59 stable beams fills 2911 Initial luminosity 4.5x10^33 cm-2 s-1 due to heavy losses on B2
  • Tuning of the injectors and in particular in the PS Booster with reduction of the emittance by about 10%. Increased chromaticity in ramp and squeeze by +5 units in both planes
  • 12:46 End of fill #2911. OP dump to refill with fresh beams and new machine settings. 61 pb-1 in 4.8 h.
  • Filling #2912. Slowed by some intensity fluctuations due to bad injection in the SPS. Beam quality check catching them
  • Very clean ramp and squeeze. Still dip in lifetime in collision to 2 hours.
  • 15:28 STABLE BEAMS #2912. Initial luminosity 6.4x10^33 cm-2 s-1
  • 23:12 dump by FMCM (thunderstorm)
  • thunderstorm: magnet trips in several sectors, SPS down
9:00 meeting
Friday 3 August
  • 00:22 Stable beams #2905 Initial luminosity 5.9x10E33
  • 00:40 Beam dump due to trip of RQ5.L8 ~6 pb-1 in 20 mins. SEU?
  • 03:43 Stable beams #2906. Initial luminosities 6x1033 cm-2s-1
  • 04:38 beam dump due to UFO @ MKI2 (156% of dump threshold). 18 pb-1 in 55 mins
  • 06:28 Fill #2907. Dumped once more by UFO during the first part of the squeeze (100 %!!!)
  • 7:30 injection for fill 2908. Average bunch intensity 1.53E11 at injection. Still rather large emittances, especially B2. High losses during ramp and squeeze..Soem activity in B1 and B2 not clearly correlated to instabilities (chrip was active)
  • 09:17 Stable beams #2908. Initial L ~6.0E33 cm-2s-1.
  • LHCb reached 1 fb-1 in 2012
  • 19:12 beam dump due to trip of RCBXV1.R5. 134 pb-1 in ~10 hours
  • 21:35 intervention on RCBXV1.R5 power module finished
  • 21:45 RD.34.LR3 tripping repeatedly while going to STANDBY
  • 23:11 piquet re-set the fault, start pre-cycling
8:30 meeting
Thursday 2 August
  • 12:09 dump by OP: 12.5 hours in stable beams; Integrated lumi: ~149 pb-1.
  • 13:38 problem with SPS B2 extraction
  • 15:13 Vacuum interlock on transverse damper (VB1 and HB2)
  • 16:09 Transverse damper OK but still issues with SPS extraction interlock
  • 16:23 PS main power supply down
  • 17:00 Problem with QPS controller on RQD.A23 (33.L3). Access required.
  •  19:36 End of access: power supply of the QPS detection unit replaced. In the shadow  verification of the RQF.A12 energy extraction (following spurious opening on sunday night)
  • 19:44 Lost patrol in PM18 (door YCPS01=PM18). Already happened last friday. Piquet intervention to replace relay on the door
  • 21:20 Patrol completed and machine closed - Precycling
  • 22:30 injection finally
8:30 meeting
Wednesday 1 August
  • 8:42 - 11:30 end of fill study - octupole polarity inversion
  • 12:30-14:00 Test of tune measurement with the damper
  • 15:59 Stable beams fill 2900. Initial L ~6.0E33 cm-2s-1;
  • 19:27 Beam dump due to instability (could be HW damper problem). 60 pb-1 in 3.5 h.
  • ADT B2H noise problems
  • 23:32 stable beams fill 2902. Initial L ~6.0E33 cm-2s-1;
8:30 meeting
Tuesday 31 July
  • 04:00 dump fill 2893: spurious BLM trigger in IP7
  • 6:40 stable beams fill 2894. 1 hour in stable beams; Integrated lumi: ~20 pb-1
  • 7:36 beam dump due to trip of RCBXH1.R1
  •  9:16 dump on IP6 BPMs during filling
  • 11:16 stable beams fill 2896. Initial L ~6.0E33 cm-2s-1; 10 min in stable beams; Integrated lumi: ~3 pb-1
  • 11:27 dump by trip of RQTL7.R7B1
  • Delays to inject due to: BLMs, SPS access, PS
  • 14:50 injection for fill 2897
  • 16:50 Beam dump when going in collision (B1H instability)
  • Trimmed chomaticity during ramp and squeeze according to measurements performed on sunday
  • 19:20 Stable beams fill 2898. Initial luminosity~6.0E33 cm-2s-1; ~161pb-1  in 13 h 20 min stable beams.
8:30 meeting
Monday 30 July
  • 0:14 Dump fill 2891, QPS trigger RQF.A12, integrated L ~97 pb-1.
  • 4:29 stable beams fill 2892. Initial L ~6.1E33 cm-2s-1; 10 hours in stable beams; Integrated lumi: ~138 pb-1
  • beam dump planned at 14:30 for VIP visit
  • 14:27 beam dump: RF module M2B1 crowbar event: detected arc in klystron probably due to vacuum degradation  
  • 14:30 - 20:00 access (IP5 visit, RF, Atlas, Alice, cryo, BCT, EPC) followed by patrol in IP5
  • 23:22 Stable beams fill 2893. Initial L ~6.0E33 cm-2s-1; 4.5 hours in stable beams; Integrated lumi: ~76 pb-1
8:30 meeting